The employees at a General Motors assembly plant are polled as they leave work. Each is asked, �What brand of automobile are you riding home in?� The random variable to be reported is the number of each brand mentioned. Is x a binomial random variable? Justify your answer.
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

i don't believe so.
for x to be a binomial random variable, it has to follow the rules of a binomial distribution which state that the probability of a value x is determined by the binomial distribution equation.
the solution has to be based on a probability that it is x versus a probability that it is not x.
here's a sample binomial distribution:
n = 3
p = .3
q = 1-p = .7
p(x) = nCx*p^x*q^(n-x)
The probability of y occurrences of x is determined by this formula.
you then get a distribution that approximates the normal distribution.
the greater the value of n, the closer the approximation.
that is not the case with what cars people are riding home in.
that distribution has no rhyme or reason and would give you a totally different type of distribution.
based on that i say no.
i'm reasonably confident this answer is correct, but you should double check with somebody else before committing to this answer.



Answer by ikleyn(53937)

.
The employees at a General Motors assembly plant are polled as they leave work.
Each is asked, �What brand of automobile are you riding home in?�
The random variable to be reported is the number of each brand mentioned.
Is x a binomial random variable? Justify your answer.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


If there were a mathematically literate expert supervising the work of the authors
who create problems for this forum, this supervisor would give the author of this problem
the lowest possible grade for the utter nonsense they spout.